OpenStudy (anonymous):

Perform this operation and report the answer to the correct number of significant figures: What is the mass of a cube of aluminum that is 4.0 cm on each edge? The density of aluminum is 2.7 g/cm^3. Show all calculations leading to an answer.

OpenStudy (swagster23780):

@AMYCARTER

OpenStudy (anonymous):

I know that, Vcube = length*width*height Vcube = 4.0cm*4.0cm*4.0cm Vcube = 64cm^3 2.7g/cm^3 = m/64cm^3 m = 2.7g/cm^3*64cm^3 m = 172.8g My issue is with the significant figure part.

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

The two pieces of information that the problem gives are side of the cube = 4.0cm density = 2.7g/cm^3 Both of those have 2 significant figures, and to keep the correct amount of sig figs you use the smallest number of figures (both of these are the same so no worries about chosing). All you need to do is round your answer down to 2 sig figs, and then probably present it in scientific notation.

OpenStudy (anonymous):

Would it not just be 1.7g?

OpenStudy (anonymous):

Close. It's be " 1.7x10^something," otherwise the answer differs by LOT

OpenStudy (anonymous):

1.7x10^2?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

yup yup ^_^ Significant figures can kinda be seem as rounding rules, so they shouldn't change an answer super drastically. Nice job.